OpenStudy (anonymous):

How the heck do I figure out the equations for the first derivative.. Simple explanation please because I am extremely confused

OpenStudy (anonymous):

Do you mean stuff like \[f(x) = x^n\]\[f'(x) = nx^{n-1}\]

OpenStudy (anonymous):

I don't even know what I mean I'm just lost

OpenStudy (anonymous):

The derivative gives the rate of change. Or in other words the slope.

OpenStudy (anonymous):

Oh.. I just don't get how it relates to the relative extrema like f(x)=3x^2+5x+2.. Like when you end up with 3 equations or something

OpenStudy (anonymous):

Remember \(y = mx + b\)? M is the slope. So let's write that as \(f(x) = mx^1 + b\) and take the derivative using that formula above. That's \(f'(x) = 1mx^{1-1} = m\) so, yeah, the slope. (B just goes away because it's a constant)

OpenStudy (anonymous):

That is just where the derivative changes sign. In other words the slope stops going up and starts going down (or stops going down and starts going up)

OpenStudy (anonymous):

So what would that look like with the equation I wrote? And what would the relative max be?

OpenStudy (anonymous):

Well first we have to take the derivative of \(f(x) = 3x^2 + 5x +2\). We can do that with the formula above! So do that first.

OpenStudy (anonymous):

That's what I don't know how to do lol

OpenStudy (anonymous):

I just showed you. You subtract 1 from the exponent and multiply it by the coefficient.

OpenStudy (anonymous):

Ohhhh!

OpenStudy (anonymous):

The derivative of \(2x^3\) is \(6x^2\) for example because it's \((2 \times 3)x^{3-1}\)

OpenStudy (anonymous):

So you just do that for every number then?

OpenStudy (anonymous):

Right. And get rid of anything with no x.

OpenStudy (anonymous):

(Or to be fancy about it, if \(h(x) = f(x) + g(x)\) then \(h'(x) = f'(x) + g'(x)\) )

OpenStudy (anonymous):

Okay so hold on..

OpenStudy (anonymous):

What if it's 5x? Would that stay as 5x?

OpenStudy (anonymous):

No, \(5x\) means \(5x^1\)

OpenStudy (anonymous):

So what's \((1 \times 5)x^{0}\) ?

OpenStudy (anonymous):

Isn't that the same thing? & so far I have (-3 x 2)(x^1) is that right?

OpenStudy (anonymous):

& would that be 5x?

OpenStudy (anonymous):

Was it negative? It was positive before. If so, so far so good.

OpenStudy (anonymous):

Yes it was negative.. Typo sorry

OpenStudy (anonymous):

And, no. \(x^0 = 1\). The derivative of \(5x\) is just 5. Like I showed above, the derivative of y=mx + b is just m.

OpenStudy (anonymous):

Ohhhhhh that makes sense I forgot anything to the power of 0 is 1.. So I have (-3 x 2)(x^1)+5x^1?

OpenStudy (anonymous):

You forgot to get rid of the x again. :P

OpenStudy (anonymous):

Wait.. What? :p

OpenStudy (anonymous):

It's just \(6x + 5\)

OpenStudy (anonymous):

Because it's \(6x^1 + 5x^0\)

OpenStudy (anonymous):

-6, rather.. I forgot the negative

OpenStudy (anonymous):

Ohhhhh wait! It's 5x^0 because its 5x^1-1 right!?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Multiplying times 1 and thinking of it as \(x^0\) and stuff is tedious and as you learn it you'll just drop the x but while learning it helps you understand the process

OpenStudy (anonymous):

Okayy that makes sense.. So would I write it as f(x)=-6x+5 or f'(x)=-6x+5 and what does the ' mean anyways?

OpenStudy (anonymous):

The ' is read "prime" and it just means "this is a derivative."

OpenStudy (anonymous):

And also.. How do you figure out if its a critical number? Like what does that have to do with anything?

OpenStudy (anonymous):

And ohhh

OpenStudy (anonymous):

Well, now we know the slope!

OpenStudy (anonymous):

Okay! So what now?

OpenStudy (anonymous):

Ok, well, imagine the graph of \(-3x^2 + 5x + 2\).. It's going to be an upside down parabola.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So, where on the parabola is the largest that function is ever going to get? (Not the exact value, just, where on the shape is it)

OpenStudy (anonymous):

The tip of the curve?

OpenStudy (anonymous):

Right! And what's the slope there? (imagine a tangent line)

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

No. The curve flattens out!

OpenStudy (anonymous):

So what does that mean?

OpenStudy (anonymous):

It means the slope is 0

OpenStudy (anonymous):

But how do you know that

OpenStudy (anonymous):

How do I know what?

OpenStudy (anonymous):

That it's 0.. You lost me :p

OpenStudy (anonymous):



OpenStudy (anonymous):

Here is my really bad not to scale parabola

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

The slope at any given point is the same as the slope of the tangent line

OpenStudy (anonymous):



OpenStudy (anonymous):

You're gunna hate me... What's a tangent line?:p

OpenStudy (anonymous):

It gets a little circular... but the tangent line is the line you draw next to the curve that intersects that point and has the same slope as the curve at that point.

OpenStudy (anonymous):

Would it help to try a completely different way of thinking about it?

OpenStudy (anonymous):

Probably :p

OpenStudy (anonymous):

Ok. You see how the parabola goes up and then goes down? In other words, the slope is positive, then it's negative.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well, since it's a continuous line, to go from positive to negative means it changed direction. So it had to have been 0 at some point.

OpenStudy (anonymous):

And whenever that happens, it WILL be 0?

OpenStudy (anonymous):

For the most part. For polynomial functions with smooth curves like this, definitely.

OpenStudy (anonymous):

Okay so that means its a critical number?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

& Why is that important

OpenStudy (anonymous):

You're trying to find the relative max, right?